OpenStudy (anonymous):

Prove the identity: (cosx-cosy/sinx+siny)+(sinx-siny/cosx+cosy)=0

OpenStudy (anonymous):

\[\begin{align*} LHS&=\frac{\cos x-\cos y}{\sin x+\sin y}+\frac{\sin x-\sin y}{\cos x+\cos y}\\\\ &=\cdots\times\frac{\cos x+\cos y}{\cos x+\cos y}+\cdots\times\frac{\sin x+\sin y}{\sin x+\sin y}\\\\ &=\frac{(\cos x-\cos y)(\cos x+\cos y)+(\sin x-\sin y)(\sin x+\sin y)}{(\sin x+\sin y)(\cos x+\cos y)}\\\\ &=\cdots \end{align*}\]

OpenStudy (anonymous):

So do you multiply across the top of the fraction then?

OpenStudy (anonymous):

Yes, expand, then use the identity \(\sin^2x+\cos^2x=1\). Everything will work out nicely.

OpenStudy (anonymous):

okay, thank you

OpenStudy (anonymous):

yw